OpenStudy (anonymous):

last one I promise... y=(ln(7x))/x^5.... what is y' and y''? I got y' correct. it is x^-6(1-5ln(7x))... cant figure out y''

OpenStudy (anonymous):

try rewriting and using the chain. ln(7x)x^(-5)

OpenStudy (anonymous):

first deriv.= u'v+uv' second deriv.= u''v+u'v'+u'v'+uv''

OpenStudy (anonymous):

u=ln(7x) v=x^-5 v'=-5x^-6 v''=30x^-7

OpenStudy (anonymous):

i got that derivative correct. i just cant find y''

OpenStudy (anonymous):

Im blanking on ln deriv. for some reason im thinking its x^-1

OpenStudy (anonymous):

if you find all the parts just plug it in.

OpenStudy (anonymous):

u=ln(7x) u'=7/7x u"=quotient rule?

OpenStudy (anonymous):

yes ln deriv is 1/x or x^-1

OpenStudy (anonymous):

u'=7x^-1 u''=-7x^-2

OpenStudy (anonymous):

so second deriv.= (-7x^(-2))(x^-5)+2(-5x^(-6))(7x^(-1))+ln(7x)30(x^(-7))

OpenStudy (anonymous):

clean it up and youre good

OpenStudy (anonymous):

I had just got that from your set up of u"v+u'v'+u'v'+uv"! thanks! my professor never gave me that set up!

OpenStudy (anonymous):

welcome.